Solve the system of equations step by step?
x-2y+3z=-7
4x+5y+z=4
-x+y-2z=5

Answer by scott8148(6628)

add 3rd eqn to 1st eqn ___ -y + z = -2 (A)

multiply 3rd eqn by 4, then add to 2nd eqn ___ 9y - 7z = 24 (B)

multiply eqn (A) by 7, then add to eqn (B) ___ 2y = 10

substitute back the value of y to find x and z